# LLaMIPa: An Incremental Discourse Parser

### v1

@@ Abstract
This paper provides the first discourse parsing experiments with a large language model (LLM) finetuned on corpora annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). The result is a discourse parser, LLaMIPa (LLaMA Incremental Parser), which is able to more fully exploit discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units. Furthermore, it is able to process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
@@ 1 Introduction
Understanding a conversation requires more than just grasping the contents of its utterances. One must also understand how different utterances relate to one another: whether an utterance answers a question or serves to explain or correct something else that was said, for example. Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges or separate nodes.
@@
Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inference required to predict the graph elements.
@@
Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2018), which offer richer representations of linguistic content, and can address factor (ii). But given their limited input size, they are confined to look at one small part of a discourse at a time and so cannot really address factor (iii).
@@
Generative large language models (LLMs) can substantially improve discourse parsing. They not only encode large amounts of lexical and conceptual information, addressing problem (ii) but their increased context windows permits an incremental parsing strategy to address problem (iii). Such a strategy makes sense cognitively and linguistically, as prior discourse structure clearly has an effect on inferences about which subsequent discourse units are attached (the link task) and how they are attached—i.e. what semantic relations link them (the link+relation task) Asher and Lascarides, (2003). Further, an incremental approach naturally applies to downstream applications like conversational agent, in which the information about the previously inferred discourse structure must be updated and available in real time.
@@
We develop an incremental parsing model LLaMIPa (LLaMa Incremental Parser), by fine-tuning LLaMA3, that uses previously inferred discourse structure to simultaneously predict the link and the link+relation tasks between new discourse units. We show LLaMIPa substantially outperforms the state of the art on both tasks.
@@
We demonstrate LLaMIPa’s performance using three datasets annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multi-party chats from an online game Asher et al., (2016). We further test (STAC-based) LLaMIPa on an out of domain dataset, Molweni Li et al., (2020).
@@
Not only does LLaMIPa outperform prior models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i). In particular, LLaMIPa can identify cases in which a single EDU has two parent EDUs—a situation that arises naturally in both multiparty and situated dialogue where a speaker can respond to multiple channels (conversational or multimodal) at once. Most current SDRT based parsers cannot handle such structures Shi and Huang, (2019); Liu and Chen, (2021); Wang et al., (2021); Chi and Rudnicky, (2022).
@@
In what follows, Section 2 provides background on discourse structure. Section 3 presents related work, while Section 4 details the multiparty corpora we use for training and testing. Sections 5 and 6 describe our model and results. Section 7 describes ablation experiments, and then we conclude in Section 8.
@@ 2 Background
Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999). Such relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in the discourse graph.
There are two main theories that have investigated complete discourse structures for texts: Rhetorical Structure Theory (RST; Mann and Thompson,, 1987) and SDRT (Asher,, 1993; Asher and Lascarides,, 2003). RST builds discourse trees over EDUs with the semantic relation labels on the head nodes dominating subtrees.
SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) structure in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs). EEUs represent salient aspects of the nonlinguistic context in which a conversation takes place, and may be connected to the discourse DAG. SDRT also allows subgraphs, or complex discourse units (CDUs) to act as terms of semantic relations.
Discourse parsing requires addressing three tasks. The segmentation task involves decomposing a text and (descriptions of) actions into EDUs and EEUs. The link task involves identifying pairs of EDUs and/or EEUs that are semantically related via discourse relations—that is, the arcs in the discourse graph. The link+relation task involves identifying the labels for each attached pair.
@@ 3 Related work
Traditionally with SDRT style parsers, gold segmentation is assumed Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked and then labels are subsequently predicted for these links. While we assume gold segmentation, our proposed architecture combines the attachment and labeling tasks.
Several strategies have evolved to take account of the complex information required for discourse parsing within a dependency based paradigm like that proposed by SDRT or DLTag (Forbes et al.,, 2003). Research has shown that semantic features of EDUs are crucial for solving the link+relation task Mann and Thompson, (1987); Asher and Lascarides, (2003). Older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) thus used manual features about EDU pairs and simple ML models to build a local model to predict link and link+relation label in a pipeline fashion. A local model uses information just about pairs of EDUs that are to be attached; it does not exploit information about the discourse structure around that pair. Afantenos et al., (2015); Perret et al., (2016) also added decoding mechanisms with post hoc constraints on the output structures, following Muller et al., (2012). Although exploiting little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023). We do a detailed comparison between an encoder based local parsing model and LLaMIPa.
Shi and Huang, (2019) were the first to obtain significant discourse parsing results using a deep learning architecture on STAC. They attempted to capture incremental and contextual effects in their model by training a supplemental Structured Encoder to incrementally update attachment paths (sequences of parent-child EDUs). However, Wang et al., (2021) showed that the model obtains similar scores with or without this encoder.
@@
Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and edge embeddings. The model uses a Hierachichal GRU gate for contextual EDU representations. They then apply the SSA-GNN to capture implicit structural information between EDUs, using a Structure-Aware Scaled Dot-Product Attention Zhu et al., (2019); Wang et al., (2020) to update edge and EDU representations. They further train a teacher network to supplement the standard classification loss with an auxiliary loss. Chi and Rudnicky, (2022) developed a more complex model based on Shi and Huang, (2019).
None of the models by Shi and Huang, (2019); Liu and Chen, (2021); Guz et al., (2020); Wang et al., (2021); Chi and Rudnicky, (2022) predict multiple parents for an EDU (or EEU). The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023); Thompson et al., (2024)’s model BERTLine. BERTLine uses BERT embeddings for EDUs and EEUs, along with a linear layer for features like speaker change and distance information, to build a local model and predicts link and link+relation in pipeline fashion. However the relation labeling uses a multitask architecture to capture informational dependencies between link and relation labeling decisions. Bennis et al., (2023) show that their model is competitive with approaches that try to build in more contextual knowledge Liu and Chen, (2021); Wang et al., (2021). Thompson et al., (2024) uses BERTLine on the MSDC, which makes it a natural baseline for our work on situated, multiparty conversation.
Recent work has investigated using LLMs to predict discourse structure, by finetuning LLaMA2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024). The reported results on RST corpora are not directly comparable to ours, but they are significantly lower numerically. Chan et al., (2024); Yung et al., (2024); Fan et al., (2024); Henning and Stede, (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations. The results are significantly worse than dedicated discourse parsers; testing with 3-shot learning on a version of the STAC corpus (STAC-L; see Section 4), Chan et al., (2024) reports an F1 score of 21.3 for link prediction and 7.4 for relation prediction by ChatGPT.
@@ 4 Data sets
MSDC. The MSDC provides full discourse structures in the style of SDRT to dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019). In the MDC, dialogues involve a builder, whose task is to build block formations in a simplified Minecraft world, and an architect, who guides the builder by giving instructions through chat. The MSDC annotates discourse relations between both chat moves (EDUs) and linguistic descriptions of the builder’s pick and place actions (EEUs), provided by the MDC logs.
The MSDC’s multimodal nature presents many opportunities to simultaneously respond to a conversational move and a building move, and indeed, the MSDC has numerous discourse units with multiple parents (MPDUs). Figure 1 shows an excerpt from the MSDC with two examples. For instance, after placing two blocks, the builder asks “like that?”. The architect’s reply, “not quite,” not only serves to answer the builder’s question but simultaneously to correct the builder’s last moves. The following two elaborating discourse moves fill out the full content of the Correction.
Figure 1 shows another important aspect of the MSDC, namely that builders tend to perform sequences of pick and place moves uninterrupted by chat moves. These action sequences are grouped together as CDUs in the MSDC, indicated by dashed boxes in Figure 1. Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent them from drowning out the more complex and interesting linguistic interactions in the corpus. We used the MSDC with this modification.
@@
STAC contains MPDUs due to both multimodal interactions and multiparty discourse, in which one player can simultneously respond to multiple players at once. Relevant statistics for STAC and the MSDC are found in Table 1.
STAC-Sit, like the MSDC, also contains sequences of game-only moves. Bennis et al., (2023) compressed such sequences into individual EEUs and we follow this choice. Unlike the MSDC, however, STAC contains numerous CDUs for chat moves and for chat/game interactions (as opposed to CDUs of action sequences only). Following Bennis et al., (2023) and most of the literature on SDRT parsing, we did not treat CDUs directly but used the “flattening" procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs. A final simplication of the original STAC corpus involved removing EEUs that neither interact with EDUs n or figure in EEU sequences that interact with EDUs. This choice, like the one to compress EEU sequences, was made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
Molweni. The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015) a collection of exchanges from an online forum for discussing issues related to Ubuntu. Unlike the MSDC and STAC, Molweni does not contain CDUs, EEUs, or MPDUs. We used the version of Molweni Test re-annotated by Li et al., (2024).
@@ 5 LLaMIPa: An Incremental Parser
LLaMIPa leverages the sequence to sequence architecture and the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy. As it generates a discourse graph for a conversation , the LLM should base its link and relation predictions for an EDU in on the graph it has already constructed for EDUs preceding .
Formally, let be a conversation consisting of a sequence of EDUs or EEUs. Our LLM takes as input a set of of discourse and (possibly) event units and a discourse graph . At step 0, is , the head EDU, and , the null discourse graph. After that, is a set , where 1 and where is the window size, set to ; and contains links for . At each step, the model predicts a graph that extends by integrating link+relation predictions for all of the EDUs in the next dialogue turn or, if the next move is an action sequence, the EEU representing this sequence. This process repeats until LLaMIPa arrives at the final unit of , effectively creating a “sliding context window” that will eventually slide over an entire dialogue, permitting the model to potentially exploit a large amount of contextual information.
We represent relation instances as EDU index pairs connected by a relation type, e.g. RES(0,1), for the instance of a Result relation holding between EDUs 0 and 1. During training, each subsequent instance in a dialogue adds to the context the previous turn EDUs and EEUs, but we take the gold, rather than predicted, discourse structure between those units to make our input graph . is specified as a list of the relations between these units in the same format as the predicted relations.
We used a sliding context widow of 15 EDUs to optimize the number of training samples while capturing the most number of discourse links that have statistical support. As discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than 15 in the SDRT corpora.
Using our corpora formatted for incremental relation prediction, we finetuned LLaMA-2 Touvron et al., (2023), and LLaMA-3222https://llama.meta.com/llama3/ to produce LLaMIPa2 and LLaMIPa3. In order to do this without training the full model weights, we use QLoRA Dettmers et al., (2024) to train for 3 epochs. The model parameter details are in Table 6 in Appendix A. The finetuned models will be made publicly available.
Once LLaMIPa is finetuned on the training corpus, the generation procedure uses predicted discourse structure instead of the gold structures during inference (See Figure 2). Proceeding turn by turn, LLaMIPa produces a complete discourse structure for a dialogue in which both links and relations are predicted. The LLaMIPa3 model and generation code is available here333https://huggingface.co/Kqte/LLaMIPa.

### v2

@@ Abstract
This paper provides the first discourse parsing experiments with a large language model (LLM) finetuned on corpora annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). The result is a discourse parser, Llamipa (Llama Incremental Parser), that leverages discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units. Furthermore, it can process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
@@ 1 Introduction
Understanding a conversation requires more than just grasping the contents of its utterances. One must also understand how different utterances relate to one another, for example whether an utterance answers a question or serves to explain or correct something previously said. Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly, clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges.
@@
Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inferences required to predict the graph elements.
@@
Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2019); Bennis et al., (2023) which offer richer representations of linguistic content, and can address factor (ii). But given their limited input size, they are confined to look at one small part of a discourse at a time, and so cannot really address factor (iii).
@@
Generative large language models (LLMs) can substantially improve discourse parsing. They not only encode large amounts of lexical and conceptual information, addressing problem (ii), but their increased context windows permit an incremental parsing strategy to address problem (iii). Such a strategy makes sense cognitively and linguistically, as prior discourse structure has an effect on inferences about where subsequent discourse units are attached (the link task), and how they are attached, i.e., which types of semantic relations link them (the link+relation task) Asher and Lascarides, (2003). Further, an incremental approach naturally applies to downstream applications such as conversational agents, in which the information about the previously inferred discourse structure must be available and updated in real time.
@@
We develop an incremental parsing model by fine-tuning Llama3 Dubey et al., (2024). Llamipa (Llama Incremental Parser) uses previously inferred discourse structure to simultaneously predict the links and relation types between new discourse units. We show that Llamipa substantially outperforms the state-of-the-art on both the link and link+relation tasks.
@@
We demonstrate Llamipa’s performance using three datasets annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we also show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multiparty chats from an online game Asher et al., (2016). We further test Llamipa fine-tuned on STAC on an out of domain dataset, Molweni Li et al., (2020).
@@
Not only does Llamipa outperform previous discourse parsing models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i). In particular, Llamipa can identify cases in which a single EDU has two parent EDUs, a situation that arises naturally in both multiparty and situated dialogue, where a speaker can respond to multiple channels (conversational or multimodal) at once. Most current SDRT based parsers cannot handle such structures Shi and Huang, (2019); Liu and Chen, (2021); Wang et al., (2021); Chi and Rudnicky, (2022).
@@
In what follows, Section 2 provides background on discourse structure. Section 3 presents related work, while Section 4 details the multiparty corpora we use for training and testing. Sections 5 and 6 describe our model and results. Section 7 describes our ablation experiments, and we conclude in Section 8.
@@ 2 Background
Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from discourse units or EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999). These relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in a discourse graph.
There are two main theories that have investigated complete discourse structures for texts: Rhetorical Structure Theory (RST; Mann and Thompson,, 1987) and SDRT (Asher,, 1993; Asher and Lascarides,, 2003). RST assigns relations between contiguous EDUs which are connected hierarchically to further EDUs in a treelike structure.
SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs). EEUs represent salient aspects of the nonlinguistic context in which a conversation takes place, and may be connected to the discourse DAG. SDRT also allows subgraphs, or complex discourse units (CDUs) to act as terms of semantic relations. SDRT is the only framework to date that handles multiparty conversation.
Whatever the underlying paradigm, discourse parsing requires addressing three tasks. First, the segmentation task, which involves decomposing a text and (descriptions of) actions into EDUs and EEUs. The link task involves identifying pairs of EDUs and/or EEUs that are semantically related via discourse relations—that is, the arcs in the discourse graph. Finally, the link+relation task involves identifying the labels for each attached pair.
@@ 3 Related work
Traditionally, SDRT style parsers assume gold discourse unit segmentation Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked, and then labels are predicted for those links. Our proposed architecture assumes gold discourse unit segmentation, but combines the attachment and labeling tasks.
Several strategies have evolved to harness the complex information required for discourse parsing within the dependency based paradigm proposed by SDRT or DLTag (Forbes et al.,, 2003). Research has shown that semantic features of EDUs are crucial for solving the link+relation task Mann and Thompson, (1987); Asher and Lascarides, (2003). Thus, older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) input manually designed features of the EDU pairs to simple ML models to predict link and link+relation label successively. These local models use information local to the pairs of EDUs that are to be attached, but do not exploit information about the surrounding discourse structure. In order to leverage some more global information, Afantenos et al., (2015); Perret et al., (2016) apply special decoding mechanisms with post hoc constraints on the output EDU pairs, following Muller et al., (2012). Although local models exploit little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023). We do a detailed comparison between an encoder based local parsing model and Llamipa.
Shi and Huang, (2019) were the first to obtain competitive discourse parsing results using a deep learning architecture on STAC. They attempted to capture incremental and contextual effects in their model by training a supplemental Structured Encoder to incrementally update attachment paths (sequences of parent-child EDUs). However, Wang et al., (2021) show that the model obtains similar scores with or without this encoder.
@@
Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and relation embeddings. The model uses a Hierachichal GRU gate for contextual EDU representations. They then apply the SSA-GNN to capture implicit structural information between EDUs, using a Structure-Aware Scaled Dot-Product Attention Zhu et al., (2019); Wang et al., (2020) to update edge and EDU representations. They further train a teacher network to supplement the standard classification loss with an auxiliary loss. Chi and Rudnicky, (2022) developed a more complex model based on Shi and Huang, (2019).
None of the models by Shi and Huang, (2019); Liu and Chen, (2021); Guz et al., (2020); Wang et al., (2021); Chi and Rudnicky, (2022) predict multiple parents for an EDU (or EEU). The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023)’s model BERTLine. BERTLine uses BERT embeddings for EDUs and EEUs, along with a linear layer for features like speaker change and distance information, to build a local model and predicts link and link+relation in pipeline fashion. However the relation labeling uses a multitask architecture to capture informational dependencies between link and relation labeling decisions. Bennis et al., (2023) show that their model is competitive with approaches that try to build in more contextual knowledge Liu and Chen, (2021); Wang et al., (2021). Thompson et al., (2024) uses BERTLine on the MSDC, which makes it a natural baseline for our work on situated, multiparty conversation.
Recent work investigates using LLMs to predict discourse structure by finetuning Llama2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024). The reported results show improvements over the state-of-the-art for RST corpora analogous to those we find using Llamipa on SDRT corpora. Chan et al., (2024); Yung et al., (2024); Fan et al., (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations. These models fare worse than dedicated discourse parsers: Chan et al., (2024) test with 3-shot learning on a version of the STAC corpus (see Section 4 for more corpora details) and report a near 50 point drop in F1 for link and link+relation prediction compared to Shi and Huang, (2019).
@@ 4 Data sets
MSDC. The MSDC provides full discourse structures in the style of SDRT for dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019).
The MSDC represents conversational moves as well as builder move sequences, individuated into block placements and removals. Figure 1 shows an excerpt from an MSDC dialogue containing a few examples of discourse units with multiple parents, or MPDUs. In one such example, after placing two blue blocks, the builder asks, “like that?”. The architect’s reply, “not quite,” not only serves to answer the builder’s question, but also corrects the builder’s last moves. The EDU for “like that?” is an MPDU that is semantically linked to both linguistic and nonlinguistic units. Such MPDUs are an important feature of the MSDC, as they facilitate the interplay of multimodal contributions in a single discourse structure.
Another noteworthy feature of the MSDC is how nonlinguistic actions are annotated. Builder moves are usually prompted by architect instructions, but often transpire in longer sequences of block placements and removals, which sometimes overlap with other linguistic chat moves. These action sequences are grouped together in single complex discourse units (CDUs) in the MSDC, indicated by the dashed boxes in Figure 1. Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent the repetitive connections between them from drowning out the more complex and interesting linguistic interactions in the corpus. We used the MSDC with this modification.
@@
Like MSDC, STAC contains MPDUs due to multimodal interactions, but also features MPDUs which arise from multiparty interaction, in which one player can simultaneously respond to multiple players at once. Relevant statistics for STAC and the MSDC are found in Table 1.
STAC-Sit, like the MSDC, also contains sequences of moves containing only EEUs. Bennis et al., (2023) compressed these sequences into individual EEUs and we follow this choice. Unlike the MSDC, however, STAC contains numerous CDUs for chat moves and for chat/game interactions (as opposed to CDUs of action sequences only). Following Bennis et al., (2023) and most of the literature on SDRT parsing, we do not treat CDUs directly, but use the “flattening” procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs. A final simplification of the original STAC corpus involved removing EEUs that neither interact with EDUs nor figure in EEU sequences that interact with EDUs. This choice, like the one to compress EEU sequences, is made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
Molweni. The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015), a collection of exchanges from an online forum for discussing issues related to Ubuntu. Unlike the MSDC and STAC, Molweni does not contain CDUs, EEUs, or MPDUs. We use the Molweni-clean test set, a version of Molweni re-annotated by Li et al., (2024).
@@ 5 Llamipa: An Incremental Parser
Llamipa leverages the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy. As it generates a discourse graph for a conversation, the LLM bases its link and relation predictions for an elementary unit based on the graph it has already constructed for units preceding .
Formally, let be a sequence where denotes an EDU or EEU. Our LLM takes as input a set , containing EDUs and EEUs, and a discourse graph . At step , is , the head EDU, and , the null discourse graph. After that, is a set , where and where is the window size, and contains links connecting . At each step, the model extends the previous discourse graph by integrating link and relation predictions for all of the EDUs or EEUs in the current dialogue turn.
In the data, we represent relation instances as elementary unit index pairs connected by a relation type, e.g. RES(0,1), for the instance of a Result relation holding between units and . During training, each sample contains the previous EDUs and EEUs within the window size, the gold structure, which is a list of relation instances between these elementary units, and finally, the units of the current turn. The model is trained to predict the gold relation instances that connect the units of the current turn with previous ones.
As LLMs can handle longer input sequences, we use a window size which allows the model to exploit a large amount of contextual information. Since discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than in the SDRT corpora.
We finetune Llama 2 Touvron et al., (2023), and Llama 3 Dubey et al., (2024) to produce Llamipa2 and Llamipa3 using QLoRA Dettmers et al., (2023). Both the models are trained for 3 epochs. The model parameter details are in Table 6 in Appendix A. We release Llamipa3 publicly through the Hugging Face platform222https://huggingface.co/linagora/Llamipa.
We test our models in two settings. In the first setting, the models predict the relation instances for samples containing the gold discourse structure. In the second, the model uses its predicted relation instances instead of the gold structure in the input sample. That is, at each step, the structure fed to the model is just the accumulated relation instances predicted within the window up until that step (see Figure 2). This incremental setting provides a clear idea of the model’s ability to build a complete dialogue structure from scratch.